OpenStudy (anonymous):

Find an equation of the line that passes through the point (9,2) and is perpendicular to the line 2x + 3y + 5 = 0.

OpenStudy (anonymous):

slope of line given line =-2/3 so slope of required line will be =3/2 use point slope form m(x-x0)=y-y0 3/2(x-9)=y-2

OpenStudy (anonymous):

 but it is perpendicular the t its slope will be 3/2 then;  c=-23/2 now sub c in y=mx+c